a set of data is normally distributed with a mean of 1000 and standard deviation of 100.
what would be the standard score for a score of 933?
what percentage of scores is between 1000 and 933?
what would be the percentile rank for a score of 933?

Answer by stanbon(75887)

a set of data is normally distributed with a mean of 1000 and standard deviation of 100.
what would be the standard score for a score of 933?
----
z(933) = (933-1000)/100 = -67/100 = -0.67
-------------------------------------------------------
what percentage of scores is between 1000 and 933?
z(1000 + 0
So, P(933< = x <=1000) = P(-0.67<= z <=0) = 0.2486
-------------------------------------------------------
what would be the percentile rank for a score of 933?
Percentile rank = 24%ile
===============================
Cheers,
Stan H.
===============================